What are a pregnant woman's nutrient needs?

(Answer must be minimum of 150 words in length. And no copy & paste from Google please. Thx!)
Answer #1

For a pregnant woman nutrients need are as follows.....

Protein requirements 1gm/kg of body weight(referance weight of a women is 55 kg) then= 55gm daily requires for a women

Energy requirements for sedentary, moderate worker and hard/heavy workers are=1900kcal,2230kcal and 2850 kcal respectively.

If women are pregnant some additional requirements like

Energy requirements = +350kcal

(for moderate working women =2230+350= 2580kcal)

Protein requirements are= +23gm extra (like 55+23=78gm)

RDA of Fat in pregnant women =30gm/dl(Normally =20)

RDA of vit A for pregnant women =800 micrograms(N=600ug)

For Vit D = 200IU/day

Calcium = 1000 mg/day

Iron= 40mg/day etc

If you have any query we will discuss in comment section otherwise please like Thanks